OpenStudy (anonymous):

caculate the hydroxide and hydronium ions concentration? 6.9x 10 to the -3 power M HNO3?

OpenStudy (anonymous):

@UselessMass ?? NEED HELP

OpenStudy (anonymous):

@aaronq NEED HELP??

OpenStudy (anonymous):

HNO3 is one of the strong acids -log(6.9x10^-3) = 2.16 pH pOH = 14 - pH = 14 - 2.16 = 11.83 pOH To find the concentration of hydroxide the equation is 10^(-pOH) so [OH-] = 10^(-11.83) = 1.45 x 10^-12 M and concentration of hydronium was already given [H+] = [H3O+] = 6.9 x 10^-3 M